OpenStudy (anonymous):

Find the slope of the tangent line f(x)=x^2 when x=1

OpenStudy (anonymous):

got your derivative?

OpenStudy (anonymous):

calculus?

OpenStudy (anonymous):

derivative at x=1 is the slope of the tangent at x=1

OpenStudy (anonymous):

then using point-slope form \[y-y1=m(x-x1)\] you'll get the tangent line at that point

OpenStudy (anonymous):

in order to find y, you just plug f(1)

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

did you get the answer?

OpenStudy (anonymous):

\[y'=2x\] \[f'(1)=2(1)=2=m\] \[f(1)=(1)^2=1=y1\] \[y-1=2(x-1)\] \[y=2x-1\]

OpenStudy (anonymous):

The slope is 2 :)

OpenStudy (anonymous):

right...